# Anomalous Ionization in the Central Molecular Zone by sub-GeV Dark Matter

### v1

@@ II Ionization rate from sub-GeV DM annihilation in the CMZ
where $\langle\sigma v\rangle$ is the DM thermally averaged annihilation cross section, $m_{\chi}$ is the DM mass, $\rho_{\chi}$ is the DM density in the CMZ and $\frac{dN_{e}^{\textrm{ann}}}{dE_{e}}=\delta(E+m_{e}-m_{\chi})$.
@@
Numerical calculation — In our main approach, we compute the steady-state diffuse $e^{\pm}$ spatial distribution and energy spectra with a recent version De La Torre Luque (2023) of the DRAGON2 code Evoli et al. (2017, 2018), a dedicated CR propagation code designed to numerically solve the full diffusion-advection-reacceleration-loss equation for the transport of charged particles in the Galactic environment Ginzburg and Syrovatskii (1969). We use the same prescription and propagation parameters as used in Refs. De la Torre Luque et al. (2023b, c, d, a), where we model the fluxes of sub-GeV positrons from DM and other exotic positron sources using a set of propagation parameters obtained through a combined fit to CR measurements by AMS-02 Aguilar et al. (2021, 2018, 2013) and Voyager-1 Webber et al. (2002); Stone et al. (2013).We refer to Ref. De la Torre Luque et al. (2023d) for full details.
@@
The advantage of this approach is that it is able to account for the effects of diffusion and reacceleration, but it has two main caveats: the first is that it depends on the extrapolation of the propagation parameters obtained from analyses of local CR data at GeV energies (see Ref. De la Torre Luque et al. (2023d) for detailed discussion), and the second is that the code does not consider how the dynamics of diffusion can change when crossing the border of the CMZ. Regarding the latter, Ref. Phan et al. (2023) showed that crossing a very dense medium can lead to a reduction of the flux of particles within molecular clouds, thereby reducing the ionization rate. This effect is more pronounced for lower positron energies and hence lower DM masses since they produce lower energy positrons. However, we note that this effect is not so relevant in our case, given that the electrons and positrons producing ionization come mostly from the interior of the CMZ, and that even with high suppression of the $e^{\pm}$ flux the predicted ionization rate can exceed that observed at the CMZ by orders of magnitude, we will be discussed further below. Therefore, differences in the predicted ionization rates from both approaches are expected, especially at masses above a few tens of MeV where the analytical approximation is not valid anymore because the diffusion rate starts to become dominant (see Fig. 3).
III Results
The main constraints on sub-GeV DM arise from cosmological observations of the early Universe such as the effective number of relativistic species $N_{\textrm{eff}}$ and the requirement for successful Big Bang Nucleosynthesis (BBN). These observations can robustly rule out DM masses below around $1$ MeV Boehm et al. (2012, 2013); Nollett and Steigman (2015), and even slightly above, depending on certain assumptions. However, recent work such as Refs. Escudero (2019); Sabti et al. (2020), has shown compatibility with observation for DM masses above a few MeV, as long as the dominant early Universe channel for DM annihilation is into neutrinos. For higher masses, there are strong astrophysical and cosmological constraints (see a recent review Ref. Balan et al. (2024b) or a more comprehensive revew of DM candidates Ref. Cirelli et al. (2024)). The strongest constraints so far come from XMM-Newton measurements Cirelli et al. (2023b); De la Torre Luque et al. (2023d) of the X-ray Galactic diffuse emission, impact of the $e^{\pm}$ injection from DM annihilation on the CMB anisotropies Slatyer (2016); Lopez-Honorez et al. (2013), and the longitude profile of the $511$ keV line De la Torre Luque et al. (2023a).
The estimated maximum ionization rate from sub-GeV DM with mass and cross-section combinations that are allowed by CMB constraints (taken from the bottom left panel of Fig. 4 of Ref. Slatyer (2016), which are obtained from the impact of DM annihilation on CMB anisotropies) is shown in Fig. 1, where we represent in the $y$-axis the predicted ionization rate from sub-GeV DM in the CMZ for generalized NFW profiles with a slope ranging from $\gamma=1$ (standard NFW distribution), to $\gamma=1.5$ (a Moore profile Moore et al. (1999)), including what is known as a contracted-NFW (cNFW) profile, which corresponds to a $\gamma=1.25$, which is renown for being the best-fit profile for the GeV Fermi-LAT Galactic Center Excess (GCE) Ackermann et al. (2017); Di Mauro (2021). In this figure, the shaded region represents the band of ionization rate values measured at the CMZ. Dashed lines represent the predicted ionization rate for the analytical approximation, while the solid lines represent the result obtained with the numerical calculation using DRAGON2. The latter should be more representative, and we have checked that uncertainties from the diffusion parameters can affect the benchmark prediction by at most a factor of a few (see Ref. De la Torre Luque et al. (2023d)). As a consequence, our result shows that there is a large region of parameter space (of DM profile slopes and DM masses) that can explain the observed ionization rate without violating cosmological constraints.
We notice that both approaches used to compute the ionization rate are approximate, given the difficulties of treating diffusion in the CMZ, however, the parameter space accessible to explain the measured ionization rate is so large that our conclusion should remain robust. To provide some quantitative examples, consider a $\simeq 10$ MeV DM particle annihilating at a rate of $\langle\sigma v\rangle\simeq 5\times 10^{-30}$, this can produce an ionization rate of $\zeta\simeq 2\times 10^{-14}$ s-1 for a slope of $\gamma=1.25$. For the same slope, a $1$ MeV DM particle could produce the observed ionization rate for $\langle\sigma v\rangle\simeq 2\times 10^{-33}$. For a Moore profile ($\gamma=1.5$) this ionization rate is achieved for a $\langle\sigma v\rangle\simeq 1.25\times 10^{-35}$ and $m_{\chi}=10$ MeV. For ease of comparison, we show in Fig. 2 (right axis – dashed lines) the expected ionization rate for a fixed annihilation cross section of $\langle\sigma v\rangle=10^{-35}$ cm3/s, with the numerical diffusion calculation (shown also in a single plot in Fig. S2 of the SM).
In terms of what concerns constraints from Galactic probes, we remark that X-ray constraints are competitive with those from CMB, although for DM particles with masses below $\sim 10$-$20$ MeV, diffuse reacceleration of the $e^{\pm}$ pairs injected by DM De la Torre Luque et al. (2023d); Seo and Ptuskin (1994), induces uncertainties at MeV energies. We have tested the X-ray emission produced by the DM-induced $e^{\pm}$ population using the Hermes code Dundovic et al. (2021) (as in Ref. De la Torre Luque et al. (2023d)), with the same propagation parameters as used for the estimations of the ionization rate in the CMZ, finding that XMM-Newton observations cannot rule out DM with masses below a few tens of MeV, as shown in the left panel of Fig. 5. We have also calculated the bremsstrahlung emission (using the Hermes code as in Ref. Luque et al. (2024)) at the MeV scale. Although bremsstrahlung emission is expected to be more important for low DM masses Cirelli et al. (2013) and dominates over the inverse Compton emission for DM masses below a few tens of MeV Bartels et al. (2017), we find that this emission is not expected to be significant for the annihilation strength needed to explain the CMZ ionization rate and it does not represent an important constraint (see right panel of Fig. 5). As expected, the bremsstrahlung emission is not competitive with observables such as $511$ keV line, X-rays or the CMB.
IV Possible connection with the 511 keV line emission
We now investigate the tantalizing possibility that the $511$ keV line measurements around the GC are correlated with the hydrogen ionization rate in the CMZ. Our discussion is motivated by the following:
ii) Ref. De la Torre Luque et al. (2023a) showed that, due to the propagation of positrons, the longitude profile of the $511$ keV line emission can be reproduced only for DM profiles cuspier than NFW, also favouring low DM masses.
iii) The power density from DM annihilations (Eq. 5) in the CMZ is compatible with observations of the line emission at the GC for a cross-section $\langle\sigma v\rangle\lesssim 10^{-32}$ cm3/s. For example, a $1$ MeV particle in a Moore DM profile can produce $\sim 10^{52}\leavevmode\nobreak\ \rm erg/s\sim 5\cdot 10^{50}e^{+}$/yr for $\langle\sigma v\rangle\lesssim 10^{-33}$ cm3/s.
$\frac{dE}{dt}\sim\frac{1}{2}\frac{\rho_{DM}^{2}}{m_{DM}}\langle\sigma v\rangle\,,$ (5)
To test if both can be connected, we compute the $511$ keV emission in the bulge (integrated in the inner $8^{\circ}$) from the annihilation of sub-GeV DM for the different DM profiles considered in this work, following the same approach as in Ref. De la Torre Luque et al. (2024), which is consistent with previous evaluations Calore et al. (2021, 2022); De la Torre Luque et al. (2023b, c); Carenza et al. (2024). Adopting an annihilation rate of $\langle\sigma v\rangle=10^{-35}$ cm3/s, we depict the predicted $511$ keV emission in Fig. 2 (left axis – solid lines), where we include a shaded gray band that accounts for the observed emission at the bulge Siegert et al. (2016), to which we add a conservative $50\%$ uncertainty (to account for underestimated systematic uncertainties). From a comparison to the ionization rate at the CMZ (right axis of the figure – dashed lines), we observe that the annihilation rate needed to explain the $511$ line emission at the bulge is lower than the annihilation rate that explains the CMZ ionization rate, by at least a factor of a few, meaning that a simultaneous explanation of both phenomena seems difficult under the DM hypothesis. This may indicate that the source of the CMZ ionization rate is not symmetrically injecting of electrons and positrons, and could instead be dominated by ionizing protons (or only electrons).
However, we note here that while uncertainties from propagation parameters similarly affect both predictions, the estimated $511$ keV emission is also affected by the conditions for positronium production. The formation of positronium from thermal positrons largely depends on the composition, temperature and density of the medium Guessoum et al. (1991), hindering our ability to make a robust prediction of this quantity (especially around the GC). Therefore, it would be difficult to firmly rule out the DM hypothesis of the CMZ ionization rate from the estimated $511$ keV line bulge emission given that uncertainties in this estimate are larger than an order of magnitude and that the cross-sections needed to match both observations can be close (e.g. for the cNFW profile case with 1 MeV DM where the 511 keV emission is compliant at $\phi^{511\,\textrm{keV}}_{8^{\circ}}\gtrsim 10^{-3}$ cm-2 s-1 and $\zeta\simeq 10^{-16}\,$s-1, only a factor of a few below the observed rate shaded red). In addition, we also remark that there are other mechanisms that could boost the ionization rate without affecting the $511$ keV production: invoking more exotic DM scenarios (that have already been proposed in the literature Farzan and Rajaee (2020); Feng and Kang (2013); Sheng et al. (2024)), it is possible to have asymmetric DM products, i.e. a different amount of electrons and positrons produced. Another intriguing possibility that may make both anomalies compatible with DM is that the direct interaction of DM particles with the H2 molecules also produces ionization.
Figure 2: Left axis: Predicted $511$ keV bulge emission associated with the $e^{\pm}$ population from annihilations for alternative DM density profiles (solid lines). The shaded band shows the constraint set by the observed emission from the bulge, with conservative $50\%$ uncertainties. Right axis: Expected ionization rate in the CMZ from annihilation of sub-GeV DM particles into $e^{+}e^{-}$, as obtained with the numerical calculation (dashed lines). The red band illustrates the range of observed values of ionization rate in the CMZ.
Since correlations between the $511$ keV line emission and the ionization rate in the CMZ can only happen if the source is injecting simultaneously electrons or positrons, we remark that observations of the ionization rate in the CMZ or even in molecular clouds, could help to discern the nature of the source producing the puzzling $511$ keV flux around the GC. However, it is clear that any source producing the emission of $511$ keV photons at the GC must be also play a role in ionizing the CMZ. These correlations will be explored more thoroughly in a separate work.
Finally, one may also wonder if the GCE could explain or significantly contribute to the CMZ ionization rate. Ref. Silk (2018) already argued about the possible connection between the GCE and the ionization rate in the CMZ caused by ultracompact minihalos of self-annihilating DM. However, the current observations point to a low-energy source of electrons or positrons to explain the ionization rate at the CMZ, that could not explain the GCE, that is observed at GeV energies. This holds both if the source is DM annihilating or millisecond pulsars. To demonstrate this, we have computed the ionization rate in the CMZ that is produced by a DM particle that can fit the GCE (to be concrete, we take, from Ref. Calore et al. (2015), the $\sim 10$ GeV DM particle that reproduces the GCE in the $\tau$ channel for a DM distribution with slope of $\gamma=1.2$), which is of $\zeta\sim 10^{21}$ s-1  [^1] We also computed the ionization rate provided by a $\sim 50$ GeV DM particle annihilating into the $b\bar{b}$ channel that fits the GCE Calore et al. (2015) and found that this candidate would produce a lower ionization rate in the CMZ, roughly by a factor of $2$-$3$, which is likely due to the higher DM mass (therefore, lower energy products).. In addition, we have computed the ionization rate produced by a population of millisecond pulsars distributed following a NFW distribution with slope of $\gamma=1.2$, whose emission of $e^{\pm}$ is a power-law with spectral index of $2$, cut-off at $3$ GeV and total power of $10^{37}$ erg/s (that could be adequate to explain the high energy tale of the GCE, as computed in Ref. Petrović et al. (2015)). In this case, the ionization rate is also largely insufficient to explain the measurements of ionization at the CMZ: $\zeta\sim 10^{23}$ s-1. Therefore, it seems that any of the currently proposed sources of the GCE would produce a negligible ionization rate in the CMZ, even lower than that from the diffuse CR sea Phan et al. (2023).
V Discussion and conclusion
In this Letter, we show that the anomalous ionization rate observed at the CMZ can be explained by DM particles of mass below a few tens of MeV annihilating into $e^{+}e^{-}$, for DM distributions with slopes slightly above that of the NFW $\gamma>1$. The thermally-averaged annihilation cross-section required to match the observed ionization rates at the CMZ is below $10^{-33}$ cm3/s for a contracted NFW distribution ($\gamma=1.25$), and can be as low as $10^{-37}$ cm3/s for a Moore DM distribution, all well below the most stringent existing upper limits. Using both analytical approximate calculations and a more detailed numerical calculation with the DRAGON2 code, we show that the required annihilation cross-sections needed to explain the CMZ ionization rate are so low that annihilation of sub-GeV DM would produce no detectable inverse Compton, bremsstrahlung or synchrotron emission. The fact that sub-GeV DM can even exceed the observed ionization rate in the CMZ without being in conflict with these probes demonstrates that this observable can be a very powerful tool for constraining sub-GeV DM. We find that the annihilation cross section needed to explain the CMZ ionization rate is above, by at least a factor of a few, that needed to explain the bulge $511$ keV emission, whilst emphasizing that the uncertainties in the 511 keV emission make it difficult to rule out DM as the source of ionization in the CMZ presently. Therefore, we determine that the $511$ keV emission can not conclusively rule out the DM explanation of the CMZ ionization rate for DM masses below a few MeV, leaving open a window to explore if both anomalies are related (even beyond the DM hypothesis). However future experiments, most notably the imminent launch of NASA’s COSI space telescope in 2027 Siegert et al. (2020), should clarify these issues. We conclude that the annihilation of sub-GeV DM that can explain the ionization rate in the CMZ could significantly contribute to an outstanding astrophysical mystery, namely the nature of the $511$ keV line emission. Finding spatial correlations between molecular clouds and positron line emission, as well as X-ray, diffuse gamma ray and radio constraints on correlated emissions may enable us to demonstrate that the source of molecular ionization must be leptonic, i.e. via injection of electrons and positrons, in deeply embedded regions of the GC.
VI Acknowledgements
We would like to thank Marco Cirelli for helpful feedback on the draft of this letter. SB is supported by the STFC under grant ST/X000753/1. SB is supported by the STFC under grant ST/X000753/1. PDL is supported by the Juan de la Cierva JDC2022-048916-I grant, funded by MCIU/AEI/10.13039/501100011033 European Union ”NextGenerationEU”/PRTR. The work of PDL is also supported by the grants PID2021-125331NB-I00 and CEX2020-001007-S, both funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. PDL also acknowledges the MultiDark Network, ref. RED2022-134411-T. This project used computing resources from the Swedish National Infrastructure for Computing (SNIC) under project No.2022/3-27 partially funded by the Swedish Research Council through grant no. 2018-05973.
Appendix A Relevant timescales
@@
Our calculations include all the relevant Galactic CR propagation processes in a broad energy range. For signals from sub-GeV DM, it is crucial to account for diffusion, reacceleration, and energy losses, that become dominant below a few tens of MeV. The most relevant cooling processes at these energies are Coulomb and ionization losses, although we include in our simulation synchrotron, bremsstrahlung, inverse Compton and catastrophic losses (annihilation) as well. A comparison of the relevant timescales is illustrated in Fig. 3, where we show a band around the benchmark diffusion time (i.e. diffusion time of particles with our benchmark diffusion setup) to indicate the difference between the optimistic and pessimistic scenarios. Moreover, cooling is much faster than diffusion. We note that the average distance travel by a positron injected at $10$ MeV (which can be calculated as $\langle l_{\text{travel}}\rangle\approx\sqrt{2D\tau}$, being $D$ the diffusion coefficient and $\tau$ the dominant timescale) is larger than $100$ pc in a $1$ cm-3 gas, without accounting for wind advection or reacceleration (that can boost $\langle l_{\text{travel}}\rangle$ to even more than $500$ pc).
@@
Figure 3: Main timescales involved in the propagation of electrons and positrons generated by sub-GeV DM particles, for a gas density of $1$ cm-3 composed by H and a $10\%$ of He. Gas densities. The magnetic field strength here is set to $5$ $\mu$G and the energy density of the interstellar radiation fields is taken from Evoli et al. (2020). The grey band indicates the difference in diffusion time in the Optimistic and Pessimistic setups, used to calculate uncertainties in the DM bounds.
Appendix B Ionization rate for fixed cross sections
For ease of comparison, we show in Fig. 4 the predicted ionization rate in the CMZ from sub-GeV DM annihilation, for different DM distributions and a fixed annhilation cross-section of $\langle\sigma v\rangle=10^{-35}$ cm3/s.

### v3

@@ Appendix A Propagation setup and calculations
where $\langle\sigma v\rangle$ is the DM thermally averaged annihilation cross-section, $m_{\chi}$ is the DM mass, $\rho_{\chi}$ is the DM density in the CMZ and $\frac{dN_{e}^{\textrm{ann}}}{dE_{e}}=\delta(E+m_{e}-m_{\chi})$.
@@
Numerical calculation — In our main approach, we compute the steady-state diffuse $e^{\pm}$ spatial distribution and energy spectra with a recent version De La Torre Luque (2023) of the DRAGON2 code Evoli et al. (2017, 2018), a dedicated CR propagation code designed to numerically solve the full diffusion-advection-reacceleration-loss equation for the transport of charged particles in the Galactic environment Ginzburg and Syrovatskii (1969). Here, we incorporate two main changes to the DRAGON2 code: the implementation of a logarithmic spatial grid that allows us to reach sub-pc spatial scales as well as the inclusion, in the neutral gas distribution, of a region resembling the CMZ, with density set to be constant at 150 cm3 and spanning the size of the CMZ. We use the same prescription and propagation parameters as used in Refs. De la Torre Luque et al. (2023b, c, d, a), where we model the fluxes of sub-GeV positrons from DM and other exotic positron sources.
We compute the steady-state diffuse distribution of electrons and positrons injected by DM by solving the diffusion-advection-convection-loss equation Ginzburg and Syrovatskii (1969); Evoli et al. (2017)
$-\nabla\cdot\left(D\vec{\nabla}f_{e}+\vec{v}_{c}f_{e}\right)-\frac{\partial}{\partial p_{e}}\left[\dot{p}_{e}f_{e}-p_{e}^{2}D_{pp}\frac{\partial}{\partial p_{e}}\left(\frac{f_{e}}{p_{e}^{2}}\right)-\frac{p_{e}}{3}\left(\vec{\nabla}\cdot\vec{v}_{c}f_{e}\right)\right]=Q_{e}\;,$ (5)
which is solved, using a Crank–Nicolson scheme Crank and Nicolson (1947), for $f_{e}\equiv\frac{dn_{e}}{dp_{e}}$, the density of $e^{\pm}$ per unit momentum at a given position. This equation includes: i) spatial diffusion with diffusion coefficient $D$, ii) momentum losses $\dot{p}_{i}$ due to interactions with the Galactic environment, iii) momentum diffusion (or reacceleration) with diffusion coefficient $D_{pp}$, iv) convection due to Galacitc winds $\vec{v}_{c}$. Here, we note that the DM density in the GC is cut at $r<2R_{\textrm{s}}$, where $R_{\textrm{s}}$ is the Schwarzschild radius of Sgr A*, the supermassive black hole in the GC Akiyama et al. (2022). This is an important physical boundary condition to prevent the source term (Eq. (4)) diverging at $r=0$. Here, we adopt a spatial grid with a bin size varying from $0.1$ pc in the inner Galaxy to $\simeq 200$ pc, increasing progressively as the radial distance from the GC is increased. We simulate electron-positron signals from DM annihilation over a range of kinetic energies from $1$ keV to a few GeV, with an energy resolution of 5%. We ensure convergence of the simulations using a variable time step (see Sect. 4.5 of Ref. Evoli et al. (2017)), with a minimum time step of $0.1$ kyr and a maximum time step of $64$ Gyr, that cover the relevant timescales involved, similar to what was done in Ref. De la Torre Luque et al. (2023a).
The diffusion coefficient employed is determined through a combined fit to CR measurements by AMS-02 Aguilar et al. (2021, 2018, 2013) and Voyager-1 Webber et al. (2002); Stone et al. (2013), and is given by $D(R)=D_{0}\beta^{\eta}\left(R/R_{0}\right)^{\delta}$, where $D_{0}=1.2\times 10^{29}$ cm2/s is its normalization at $R_{0}=4$ GV, $\beta$ is the speed of the particles in units of the speed of light, $\eta=-0.75$ and $\delta=0.49$. These are found for a halo height of $8$ kpc. We use these as our reference propagation parameters but also compare the results obtained with extreme variations in the next section (Sec. B). Unfortunately, we have no information about the exact propagation parameters within the CMZ, however, the use of extreme variations of the parameters obtained from CR analyses serves to estimate how much our results can change from our reference setup. We refer to Ref. De la Torre Luque et al. (2023d) for full details on the impact of the propagation parameters in the flux of the electrons and positrons injected by sub-GeV DM.
@@
The advantage of this approach is that it is able to account for the effects of diffusion and reacceleration, but there are two main caveats: the first is that it depends on the extrapolation of the propagation parameters obtained from analyses of local CR data at GeV energies (see Ref. De la Torre Luque et al. (2023d) for detailed discussion), and the second is that the code does not consider how the dynamics of diffusion can change when crossing the border of the CMZ. Regarding the latter, Ref. Phan et al. (2023) showed that crossing a very dense medium can lead to a reduction of the flux of particles within molecular clouds, thereby reducing the ionization rate. This effect is more pronounced for lower positron energies and hence lower DM masses since they produce lower energy positrons. However, we note that this effect is not so relevant in our case, given that the electrons and positrons producing ionization come mostly from the interior of the CMZ, and that even with high suppression of the $e^{\pm}$ flux the predicted ionization rate can exceed that observed at the CMZ by orders of magnitude, which will be discussed further below. Therefore, differences in the predicted ionization rates from both approaches are expected, especially at masses above a few tens of MeV where the analytical approximation is not valid anymore because the diffusion rate starts to become dominant (see Fig. 3 in the next section).
Appendix B Relevant timescales and propagation uncertainties
@@
Our calculations include all the relevant Galactic CR propagation processes in a broad energy range. For signals from sub-GeV DM, it is crucial to account for diffusion, reacceleration, and energy losses, that become dominant below a few tens of MeV. The most relevant cooling processes at these energies are Coulomb and ionization losses, although we include synchrotron, bremsstrahlung, inverse Compton and catastrophic losses (annihilation) in our simulation as well. A comparison of the relevant timescales is illustrated in Fig. 3, where we show a band around the benchmark diffusion time (i.e. diffusion time of particles with our benchmark diffusion setup) to indicate the difference between two extreme diffusion setups, the optimistic and pessimistic scenarios, which are explained below. Moreover, cooling is much faster than diffusion. We note that the average distance traveled a positron injected at $10$ MeV (which can be calculated as $\langle l_{\text{travel}}\rangle\approx\sqrt{2D\tau}$, being $D$ the diffusion coefficient and $\tau$ the dominant timescale) is of tens of parsecs in a $150$ cm-3 gas, without accounting for wind advection or reacceleration (that can boost $\langle l_{\text{travel}}\rangle$ to even more than hundred parsecs).
@@
Figure 3: Main timescales involved in the propagation of electrons and positrons generated by sub-GeV DM particles, for a gas density of $150$ cm-3 composed of H2 and $10\%$ of He. The magnetic field strength here is set to $150$ $\mu$G and the energy density of the interstellar radiation fields is taken from Evoli et al. (2020). The grey band indicates the difference in diffusion time in the Optimistic and Pessimistic setups, used to calculate uncertainties from diffusion in our estimations of the CMZ ionization rate.
Given that the propagation conditions around the GC are currently so challenging to evaluate, we employ the propagation parameters that reproduce the local CR observations, in particular AMS-02 and Voyager-1 data, as explained in detail in Ref. De la Torre Luque et al. (2023d). This allows us to make a reasonable estimate of the expected ionization rate in the CMZ. Similarly to what was done in Ref. De la Torre Luque et al. (2023d), in order to understand the impact of uncertainties in the propagation parameters, we have studied two opposite and extremal setups:
  • The “optimistic” setup, where CRs diffuse much faster than in our reference setup. Here, the normalization of the diffusion coefficient, $D_{0}$, is doubled (i.e. in this setup $D_{0}=20.4\times 10^{28}$ cm/s2) with respect to our reference value and have high reacceleration (which effectively contributes to a faster diffusion since it promotes low energy particles to higher energy particles), with $V_{A}=40$ km/s.
  • The “pessimistic” scenario, where diffusion is significantly reduced - to a value of $3.8\times 10^{28}$ cm/s2 - and there is no reacceleration.
As we show in the left panel of Fig. 4, the average ionization rate predicted in the CMZ changes by a factor of a few in our estimations, being more important at higher masses than at lower masses. This is because, in the MeV energy region, the dominant timescale for propagation is energy losses, which reduces the impact of uncertainties in diffusion. For completeness, we show, in the left panel of Fig. 4, uncertainties in the expected radial profile of the ionization rate (assuming an NFW DM distribution), where this becomes clearer. These uncertainties can be easily compensated for with a different value of the annihilation cross-section and provide the required ionization rate in the CMZ even for the most conservative of our estimations, since the annihilation cross-section required is largely unconstrained in this mass range.
Figure 4: Left panel: Uncertainties in the average ionization rate from DM for different DM distributions. The bands represent the different from the estimation with the Optimistic and Pessimistic setups. Right panel: Similar to the left panel but for the radial profile of the ionization, and for an NFW profile.
Appendix C Ionization rate for fixed cross-sections
For ease of comparison, we show in Fig. 5 the predicted ionization rate in the CMZ from sub-GeV DM annihilation, for different DM distributions and a fixed annihilation cross-section of $\langle\sigma v\rangle=10^{-35}$ cm3/s. As shown in the bottom panel of Fig. 1, the predicted ionization profiles become very flat when including the effects of propagation. In this figure, we show the profile of the ionization rate for different DM masses, comparing the predictions using a Moore profile and an NFW one. From this figure, one can see that for a NFW DM distribution, the radial profile of the ionization rate is quite flat, varying by no more than a factor of 3 for a 5 MeV DM particle, while much flatter at higher masses, for which the $e^{\pm}$ diffuse faster. In comparison, the ionization rate in the CMZ, measured employing different techniques, has been observed to be scattered within values that cover two orders of magnitude Ravikularaman et al. (2024). For the Moore profile, representing the cuspiest DM distribution used here, the ionization rate at different radial distances varies between a factor of a few and slightly more than an order of magnitude, which can be still compatible with the observations. Therefore, we conclude that DM is expected to produce smooth ionization profiles in the CMZ as measured. Point-like sources or Sgr A* fail on this point. However, a smooth population of sources injecting low-energy particles within the CMZ can provide a similarly good explanation.